Problem 4. The life X, in hours, of a certain device, has a pdf 100 , t 100 0, t< 100 (a) What are the probability that this device will survive 150 hours of operation? (b) Find the life expectancy of the device.
Answer #1

a)probability that device will survive 150 years =P(T>150)=150 f(t) dt =150 100/t2 dt =-100/t |infty150

=100/150=2/3 =0.6667

b)

expectancy =E(t)=100 t*f(t) dt =100 100/t dt= 100*ln(t) |infty100 =100*ln(infty)-100*ln(100) ~infty

as ln(infty) tends to infinite ; therefore life expectancy of the device does nt exist. or infinite